OpenStudy (anonymous):

The Smiths plan to recarpet their family room, which measures 15 ft. by 20 ft. How many square yards of carpet are needed?

OpenStudy (anonymous):

100/3 yds^2

OpenStudy (anonymous):

15*20=300ft convert to yard 3y 1ft ---- = ---- cross and time y? 300ft 300ft=3*300 300ft=900 yards